# Bridging Synthetic and Real Worlds for Pre-training Scene Text Detectors

### v1

@@ Abstract
Existing scene text detection methods typically rely on extensive real data for training. Due to the lack of annotated real images, recent works have attempted to exploit large-scale labeled synthetic data (LSD) for pre-training text detectors. However, a synth-to-real domain gap emerges, further limiting the performance of text detectors. Differently, in this work, we propose FreeReal, a real-domain-aligned pre-training paradigm that enables the complementary strengths of both LSD and unlabeled real data (URD). Specifically, to bridge real and synthetic worlds for pre-training, a novel glyph-based mixing mechanism (GlyphMix) is tailored for text images. GlyphMix delineates the character structures of synthetic images and embeds them as graffiti-like units onto real images. Without introducing real domain drift, GlyphMix freely yields real-world images with annotations derived from synthetic labels. Furthermore, when given free fine-grained synthetic labels, GlyphMix can effectively bridge the linguistic domain gap stemming from English-dominated LSD to URD in various languages. Without bells and whistles, FreeReal achieves average gains of 4.56%, 3.85%, 3.90%, and 1.97% in improving the performance of DBNet, PANet, PSENet, and FCENet methods, respectively, consistently outperforming previous pre-training methods by a substantial margin across four public datasets. Code will be released soon.
@@ 1 Introduction
Figure 1: We present different pre-training paradigms for enhancing text detectors. (a) represents the existing pre-training paradigm by leveraging visual and textual knowledge. (b) denotes our paradigm by leveraging the intrinsic qualities of real-world unlabeled images with annotations derived from synthetic labels. Experiments demonstrate that our method achieves comprehensive performance improvements on public datasets.
@@
In contrast, recent research efforts [34, 31, 44] have focused on pre-training scene text detectors using large-scale labeled synthetic data (LSD), which leverage both visual and textual knowledge to extract text feature representations, as illustrated in Fig. 1(a). For example, VLPT [31] formulates three cross-modality pretext tasks to enhance the performance of the backbone, and oCLIP [44] focuses on aligning visual and textual information to assist text detectors. Nevertheless, these well-curated designs are limited to using synthetic images due to their reliance on labeled word transcriptions. This limitation will lead to a noticeable domain gap while models are fine-tuned on real data, negatively impacting text detection performance.
Figure 2: We present different domain bridging ways between labeled synthetic data (LSD) and unlabeled real data (URD). (I) Mixup may lead to unexpected pixel-wise ambiguities; (II) CutMix may cause the absence of many text regions and incomplete semantics of texts; (III) ClassMix will introduce salient boundary priors; (IV) GlyphMix (ours) effectively preserves the semantic information of text without causing real domain drift and bringing extra boundary priors. Experiments demonstrate that our domain bridging method significantly performs better in aligning with the real domain, surpassing other methods by a substantial margin.
@@
Differently, given unlabeled real data (URD) readily available, our objective is to harness the complementary strengths of LSD and URD for enhancing the text detector pre-training, as shown in Fig. 1(b). The primary challenge is how to effectively modulate these URD and LSD without encountering real domain drift. In light of this, we initially investigate the adaptation capabilities of existing domain-bridging methods [55, 52, 26]. This involves training a classifier to distinguish between URD and LSD and subsequently assessing the confidence of these domain-bridging methods in aligning with the real domain. As shown in Fig. 2(c), their highest score is only 78.9%. The observation implies that these domain-bridging methods may introduce intermediate domains and disrupt the real domain.
@@
To conquer this domain drift issue, we propose FreeReal, a real-domain-aligned pre-training paradigm within a standard student-teacher framework, leveraging synthetic labels to guide text detector learning on real-domain-aligned images. Specifically, a glyph-based mixing mechanism, named GlyphMix, is tailored for text images to create real-world images with annotations derived from synthetic labels. This involves delineating the glyphs of synthetic images based on their internal structure and embedding them as graffiti-like units onto any unlabeled real images. GlyphMix offers several advantages: 1) extracting glyph structures in a cheap way without introducing additional boundary priors; 2) leveraging synthetic labels for free while significantly minimizing real domain drift (achieving a confidence score of 95.7%).
@@
As a byproduct, GlyphMix can bridge the language-to-language gap illustrated in Fig. 3, when adapting features learned from English-dominated LSD to URD in various languages. Specifically, given free character-level synthetic labels, GlyphMix enables the pre-trained model to easily learn character features, since characters are the underlying unit of any language. Subsequently, during fine-tuning, these learned character features are adaptively integrated according to the rules (e.g., word- and sentence-level) provided by available benchmarks.
@@
Building upon these insights, we break the trend of recent pre-training methods for scene text detectors (PSTD) that combine increasingly complex pretext task designs or more training modules to improve performance. Without bells and whistles, we focus mainly on exploring the potential of integrating the LSD and URD. Remarkably, despite using 90% less labeled data than recent SOTAs, our method still attains a 1.2% gain, effectively underscoring the superiority of our pre-training paradigm. As depicted in Table 1, FreeReal consistently outperforms recent SOTAs by a large margin. In summary, the main contributions are as follows:
@@
Figure 3: We illustrate differences in character, word, and sentence spacing between labeled synthetic images (a) and unlabeled real images (b). When leveraging their inherent properties to improve text detection, the divergence in spacing can confuse the pre-trained model, making it challenging to interpret text layout and semantic context. For example, we present some failure examples in (c), where the model predictions cover text areas well but yield low F-measure values. Green areas are filled by ground-truths, while red polygons indicate network predictions.
@@ 2.1 Scene Text Detection
Existing scene text detection methods delineate text regions by point-based or segmentation-based representations. The point-based methods [57, 20, 7, 40, 11] directly regress the bounding points to enclose text regions, and gradually increase the number of points as required to detect complex scenes, e.g., arbitrary texts. The segmentation-based representations regard arbitrary-shape texts by grouping pixel-level [38, 18, 17], character-level [1, 8], segment-level [21, 29] or contour-level [35, 41] segmentation results by post-processing.
2.2 Pre-training for Scene Text Detector
@@
Inspired by recent advancements in pre-training techniques [4, 3, 28, 9], the integration of text and visual cross-modal information into scene text detectors has garnered increasing attention. One representative work involves enhancing the backbone with text transcriptions. For example, Wan et al. [34] propose a pioneering work that pre-trains the backbone through an image-level text recognition task. Song et al. [31] formulate three cross-modality pretext tasks to boost better text representations derived from the backbone. Another representative work is to endow the entire detector with pre-trained vision-text knowledge. Inherited from CLIP [28], Yu et al. [50] incorporate the Image and Text encoders into the text detector to refine text regions using cross-modal visual-language priors. Despite the utilization of large-scale synthetic text images for pre-training, a persistent challenge remains—the synth-to-real domain gap continues to limit the detection performance. Additionally, they do not provide a solution for exploiting unlabeled real text images in the pre-training stage.
2.3 Semi-Supervised Learning
@@
Due to the scarcity of expensive and labor-intensive annotations, semi-supervised learning (SSL) methods are becoming increasingly important. These methods extract knowledge from a reduced amount of annotated images in a supervised manner, and from a larger amount of unlabeled images in an unsupervised way. In scenarios where both labeled and unlabeled data are from the same domain, the challenge involves leveraging the intrinsic qualities of the unlabeled data to improve model performance. In scenarios where labeled and unlabeled data belong to different domains, the challenge involves adapting a model trained on labeled data from a source domain to perform well on unlabeled data from a target domain. This situation is called unsupervised domain adaptation. Commonly, a self-training framework for SSL is employed, which encourages a model (student) to learn unlabeled images from pseudo-labels generated by the up-to-date optimized model (teacher). To avoid overfitting incorrect pseudo-labels, recent advancements have focused on pseudo-label refinement by improving the reliablity [53, 30] and pseudo-label regularization by cross-view consistency [42, 16, 33].
@@ 3 Methodology
Problem setting: Given a labeled synthetic set $\mathcal{D}_{l}={\{(\mathbf{X}_{l}^{(i)},\mathcal{B}_{l}^{(i)})\}_{i=1}^{N_{l}}}$ with $N_{l}$ samples, $\mathbf{X}_{l}^{(i)}\in\mathbb{R}^{H\times W\times 3}$ represents an image and $\mathcal{B}_{l}^{(i)}$ is its corresponding set of bounding boxes. $H$ and $W$ are the height and width of the image. Similarly, the unlabeled real set is denoted as $\mathcal{D}_{u}={\{\mathbf{X}_{u}^{(i)}\}_{i=1}^{N_{u}}}$ with $N_{u}$ samples. Our objective is to enhance the pre-training of scene text detectors by effectively leveraging the complementary strengths of both labeled synthetic data $\mathcal{D}_{l}$ and unlabeled real data $\mathcal{D}_{u}$.
3.1 Overview
We employ the segmentation-based text detector DBNet [17] as our baseline model, primarily due to its impressive real-time performance, versatility, and practical applicability. For notation simplification in the following introduction, we omit the superscript $i$ as mentioned above.
As illustrated in Fig. 4, our network pipeline consists of two principal branches: a student branch and a teacher branch. In the teacher branch, the teacher model parameterized by $\mathbf{f}_{t}$, is responsible for generating truncated-gradient logits $\mathbf{\hat{Y}}_{u}\in\mathbb{R}^{H\times W}$ from a weakly augmented unlabeled real image ${w(\mathbf{X}_{u})}$. The pseudo label $\mathbf{Y}_{u}\in\{0,1\}^{H\times W}$ is then obtained by binarizing the logit map $\mathbf{\hat{Y}}_{u}$. In the student branch, the process begins by generating a pixel-wise binary map $\mathbf{Y}_{l}\in\{0,1\}^{H\times W}$, where 1 is assigned to locations within all bounding boxes from annotation $\mathcal{B}_{l}$, and 0 otherwise. The labeled synthetic image $\mathbf{X}_{l}$ and the strongly augmented unlabeled real image $s(\mathbf{X}_{u})$ are combined to create $\mathbf{X}_{g}=\texttt{A}(s(\mathbf{X}_{u}),\mathbf{X}_{l})$, with $\texttt{A}(\cdot)$ denoting our proposed GlyphMix mechanism, which will be introduced later. Subsequently, the mixed image $\mathbf{X}_{g}$ serves as input to train the student model, under the supervision of mixed pseudo-labels $\mathbf{Y}_{g}=\texttt{A}(\mathbf{Y}_{u},\mathbf{Y}_{l})$, by minimizing loss in the following:
@@ 3.1 Overview
To filter out the unreliable predictions in mixed pseudo-labels $\mathbf{Y}_{g}$, we adaptively adjust the matrix $\mathbf{E}$ during pre-training. Specifically, for the unlabeled real image $\mathbf{X}_{u}$, we first compute the pixel-wise entropy map on $\mathbf{\hat{Y}}_{u}$ and then apply a dynamic entropy threshold $\zeta$ to obtain its positive regions $\mathbf{E}_{u}$ as follows:
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function. For the labeled synthetic image $\mathbf{X}_{l}$, the positive regions are $\mathbf{E}_{l}=\mathbf{Y}_{l}$. The matrix is subsequently formulated as $\mathbf{E}=\texttt{A}(\mathbf{E}_{u},\mathbf{E}_{l})$. In our experiments, the threshold $\zeta$ is determined by the value corresponding to the top $\gamma$ percent of the entropy map. As the reliability of pseudo-labels gradually improves, we smoothly update $\gamma$ using np.linspace(80, 20, t), where t denotes the total number of training iterations.
@@
Figure 4: Our network pipeline in a student-teacher framework.
@@ 3.2 GlyphMix
When utilizing both LSD and URD within a standard student-teacher framework, a substantial synth-to-real domain gap emerges. Although some efforts [55, 52, 26] have attempted to aggregate the advantages of synthetic and unlabeled images through versatile transformations, they inevitably introduce intermediate domains, leading to drift within the real domain. In this section, we present a novel domain-bridging way tailored for text images, allowing for the free use of synthetic labels in a cost-effective way while ensuring consistent alignment with the real domain.
1) Graffiti-like Inter-domain Mixing. Given an input synthetic image $\mathbf{X}_{l}$, its corresponding bounding box annotations $\mathcal{B}_{l}$ and converted pixel-wise binary map $\mathbf{Y}_{l}$, our goal is to perform character segmentation. This process identifies all character structure regions and produces a mask $\mathbf{M}_{G}$. Specifically, the bounding box annotations $\mathcal{B}_{l}$ are represented as $\{(p_{lt}^{i},p_{lb}^{i},p_{rt}^{i},p_{rb}^{i})\}_{i=1}^{n}$, where $p$ refers to the values of the vertice along the abscissa and ordinate, and $n$ denotes the number of bounding boxes.
@@
Given the pseudo label $\mathbf{Y}_{u}$ generated by the teacher model on an unlabeled real image $\mathbf{X}_{u}$, we use $\mathbf{M}_{G}$ to generate the inter-domain mixing image and its corresponding label. This process can be formulated as follows:
@@
where 1 is an all-one matrix and $\odot$ denotes the element-wise multiplication operation. $\mathbf{M}_{G}$ indicates pixels occupied by the glyph need to be copied from the synth domain and pasted to the real domain. Thus the mixed image closely aligns with the real domain, taking advantage of the freely available synthetic labels. Put simply, this process produces real-world images with annotations.
Figure 5: The pipeline for generating the character structures $\mathbf{M}_{G}$, including three stages: cropping, clustering, and pasting, given a synthetic image with bounding box annotations.
2) Text-balanced Intra-domain Mixing. Real-world text images inherently exhibit complexity, and the collected unlabeled data often struggles to comprehensively represent the full spectrum of real-world scenarios. Within a student-teacher framework, intra-domain mixing serves as an effective perturbation technique, enriching the diversity of real data and enhancing feature consistency and robustness.
However, when randomly generating and applying binary masks to construct intra-domain mixing images and their corresponding labels, there is a risk of missing many text regions, especially considering that text regions typically occupy only a small fraction of an image. To mitigate this challenge, our objective is to retain more text regions while still maintaining randomness.
@@
where 1 is an all-one matrix and $\odot$ represents the element-wise multiplication operation.

### v3

@@ Abstract
Existing scene text detection methods typically rely on extensive real data for training. Due to the lack of annotated real images, recent works have attempted to exploit large-scale labeled synthetic data (LSD) for pre-training text detectors. However, a synth-to-real domain gap emerges, further limiting the performance of text detectors. Differently, in this work, we propose FreeReal, a real-domain-aligned pre-training paradigm that enables the complementary strengths of both LSD and unlabeled real data (URD). Specifically, to bridge real and synthetic worlds for pre-training, a glyph-based mixing mechanism (GlyphMix) is tailored for text images. GlyphMix delineates the character structures of synthetic images and embeds them as graffiti-like units onto real images. Without introducing real domain drift, GlyphMix freely yields real-world images with partial annotations derived from synthetic labels. Furthermore, when given free fine-grained synthetic labels, GlyphMix can effectively bridge the linguistic domain gap stemming from English-dominated LSD to URD in various languages. Without bells and whistles, FreeReal achieves average gains of 1.97%, 3.90%, 3.85%, and 4.56% in improving the performance of FCENet, PSENet, PANet, and DBNet methods, respectively, consistently outperforming previous pre-training methods by a substantial margin across four public datasets. Code will be available at https://github.com/SJTU-DeepVisionLab/FreeReal.
Keywords:
Scene text detection Pre-training paradigm Domain-bridging method
00footnotetext: Corresponding author.
@@ 1 Introduction
Figure 1: We present different pre-training paradigms for enhancing text detectors.
@@
In contrast, recent research efforts [41, 37, 52] have focused on pre-training scene text detectors using large-scale labeled synthetic data (LSD), which leverage both visual and textual knowledge to extract text feature representations, as illustrated in Fig. 1(a). For example, VLPT [37] formulates three cross-modality pretext tasks to enhance the performance of the backbone, and oCLIP [52] focuses on aligning visual and textual information to assist text detectors. Nevertheless, these well-curated designs are limited to using synthetic images due to their reliance on labeled word transcriptions. This limitation will lead to a noticeable synth-to-real domain gap when models are fine-tuned on real data, negatively impacting text detection performance. Therefore, exploring the potential of unlabeled real data (URD) is of great importance as they are readily available. However, the pre-training on URD has seldom been studied. One of the few works [43] attempts to pre-train text detectors on URD via masked image modeling [12] (step 1) and on LSD via supervised learning (step 2), which will re-transfers pre-trained weights from the real domain to the synthetic domain.
@@
Differently, our objective is to jointly leverage LSD and URD by bridging their gap for enhancing the text detector pre-training, as shown in Fig. 1(b). The primary challenge is how to effectively modulate these URD and LSD without encountering real domain drift. In light of this, we initially investigate the adaptation capabilities of existing domain-bridging methods [63, 60, 30]. This involves training a classifier to distinguish between URD and LSD and subsequently assessing the confidence of these domain-bridging methods in aligning with the real domain. As shown in Fig. 2(b), their highest score is only 78.9%. The observation implies that these domain-bridging methods may introduce intermediate domains and disrupt the real domain.
@@
To conquer this domain drift issue, we propose FreeReal, a real-domain-aligned pre-training paradigm, leveraging synthetic labels to guide text detector learning on real-domain-aligned images. Specifically, a glyph-based mixing mechanism, named GlyphMix, is tailored for text images to create real-world images with annotations derived from synthetic labels. This involves delineating the glyphs of synthetic images based on their internal structure and embedding them as graffiti-like units onto any unlabeled real images. GlyphMix offers several advantages: 1) extracting glyph structures in a cost-effective way without introducing extra boundary priors; 2) leveraging synthetic labels for free while greatly minimizing real domain drift (achieving a confidence score of 95.7%).
@@
As a byproduct, FreeReal can bridge the language-to-language gap, when adapting features learned from English-dominated LSD to URD in various languages. Specifically, given free character-level synthetic labels, FreeReal enables the pre-trained model to easily learn character features, since characters are the underlying unit of any language. Subsequently, during fine-tuning, these learned character features are adaptively combined according to the rules (e.g., word- and sentence-level) provided by available benchmarks.
@@
Building upon these insights, we break the trend of recent pre-training methods for scene text detectors (PSTD) that combine increasingly complex pretext task designs or more training modules to improve performance. Without bells and whistles, we focus on exploring the potential of integrating the LSD and URD. Remarkably, despite using 90% less LSD than recent SOTAs, our method still attains a 1.2% gain, effectively underscoring the superiority of our pre-training paradigm. As shown in Table 1, FreeReal consistently surpasses recent SOTAs by a large margin. The main contributions are summarized as follows:
Figure 2: We present the domain classification task and the performances of different domain bridging ways between LSD and URD. In the text detection area, experiments demonstrate that our domain bridging method significantly performs better in aligning with the real domain. This is easy to explain: (I) Mixup may lead to unexpected pixel-wise ambiguities; (II) CutMix may cause the absence of many text regions and incomplete semantics of texts; (III) ClassMix will introduce salient boundary priors; (IV) GlyphMix (ours) effectively preserves the semantic information of text without causing real domain drift and bringing extra boundary priors.
@@ 2 Related Work
Scene Text Detection. Existing STD methods delineate text regions by regression-based or segmentation-based representations. The regression-based methods [65, 66, 22, 8, 47, 13, 32] directly regress the bounding points to enclose text regions, and gradually increase the number of points as required to detect complex scenes, e.g., arbitrary texts. The segmentation-based methods regard texts of arbitrary shapes by grouping pixel-level [46, 21, 20, 38], character-level [1, 9], segment-level [24, 35] or contour-level [42, 48, 33] segmentation results by post-processing.
@@
Pre-training for Scene Text Detector. In early research, [11, 25, 19] proposed to pre-train scene text detectors by generating synthetic data. These data engine tools follow the same paradigm: they first stack multiple off-the-shelf models to perform semantic segmentation and depth estimation analysis on background images, and then blend pre-processed texts into these images. However, the multi-stage, pre-generated methods incur large computational costs (GPU and CPU resources) and substantial latency, making them unsuitable for online training tasks where immediate data processing is essential. Inspired by recent advancements in pre-training techniques [4, 3, 34, 10], the integration of text and visual cross-modal information into scene text detectors has garnered increasing attention. One representative work involves enhancing the backbone with text transcriptions. For example, STKM [41] proposes a pioneering work that pre-trains the backbone through an image-level text recognition task. VLPT [37] formulates three cross-modality pretext tasks to boost better text representations derived from the backbone. Another representative work is to endow the entire detector with pre-trained vision-text knowledge. Inherited from CLIP [34], TCM incorporates the Image and Text encoders into the text detector to refine text regions using cross-modal visual-language priors. A persistent challenge still remains—the synth-to-real domain gap limits the detection performance.
@@
Semi-Supervised Learning. Due to the scarcity of expensive annotations, semi-supervised learning (SSL) methods are becoming increasingly important. These methods extract knowledge from a reduced amount of annotated images in a supervised manner, and from a larger amount of unlabeled images in an unsupervised way. Commonly, a self-training framework for SSL is employed, which encourages a model (student) to learn unlabeled images from pseudo-labels generated by the up-to-date optimized model (teacher). To avoid overfitting incorrect pseudo-labels, recent advancements have focused on pseudo-label refinement by improving the reliablity [62, 36] and pseudo-label regularization by cross-view consistency [50, 18, 40]. Specifically, some domain adaption works [61, 49, 5] for STD focus on minimizing domain discrepancies in the feature space, including self-training and adversarial alignment methods. The self-training methods adopt the source data to pre-train a model for generating pseudo labels on target data, then use them to retrain the model. Adversarial alignment approaches aim to align the feature distribution and learn the domain-invariant features between the source and target domains. Recently, they have been replaced by the student-teacher framework, which significantly reduces domain gaps through dual-branch knowledge transfer [54].
Figure 3: Our network pipeline in a student-teacher framework.
@@ 3 Methodology
Overview. Given a labeled synthetic set $\mathcal{D}_{l}={\{(\mathbf{X}_{l}^{(i)},\mathcal{B}_{l}^{(i)})\}_{i=1}^{N_{l}}}$ with $N_{l}$ samples, $\mathbf{X}_{l}^{(i)}\in\mathbb{R}^{H\times W\times 3}$ represents an image and $\mathcal{B}_{l}^{(i)}=\{(p_{lt}^{i},p_{lb}^{i},p_{rt}^{i},p_{rb}^{i})\}_{i=1}^{n}$ is its corresponding set of bounding boxes. $p$ refers to the values of the vertice along the abscissa and ordinate, and $n$ denotes the number of bounding boxes. $H$ and $W$ are the height and width of the image. Similarly, the unlabeled real set is denoted as $\mathcal{D}_{u}={\{\mathbf{X}_{u}^{(i)}\}_{i=1}^{N_{u}}}$ with $N_{u}$ samples. Our objective is to jointly leverage labeled synthetic data $\mathcal{D}_{l}$ and unlabeled real data $\mathcal{D}_{u}$ by bridging their gap for enhancing the pre-training of scene text detectors.
The student-teacher framework exhibits substantial efficacy due to its dual capacity for knowledge transfer between its two branches, which is typically employed to mitigate the domain gap (e.g., data and multimodal). Accordingly, we customize it for aligning real and synthetic text data, as illustrated in Fig. 3. For notation simplification in the following introduction, we omit the superscript $i$ as mentioned above.
Specifically, the segmentation-based text detector DBNet [20] is employed as our baseline model, primarily due to its impressive real-time performance, versatility, and practical applicability. In the teacher branch, the teacher model parameterized by $\mathbf{f}_{t}$, is responsible for generating truncated-gradient logits $\mathbf{\hat{Y}}_{u}\in\mathbb{R}^{H\times W}$ from a weakly augmented unlabeled real image ${w(\mathbf{X}_{u})}$. The pseudo label $\mathbf{Y}_{u}\in\{0,1\}^{H\times W}$ is then obtained by binarizing the logit map $\mathbf{\hat{Y}}_{u}$. In the student branch, the process begins by generating a pixel-wise binary map $\mathbf{Y}_{l}\in\{0,1\}^{H\times W}$, where 1 is assigned to locations within all bounding boxes from annotation $\mathcal{B}_{l}$, and 0 otherwise. The labeled synthetic image $\mathbf{X}_{l}$ and the strongly augmented unlabeled real image $s(\mathbf{X}_{u})$ are combined to create $\mathbf{X}_{g}=\texttt{A}(s(\mathbf{X}_{u}),\mathbf{X}_{l})$, with $\texttt{A}(\cdot)$ denoting our proposed GlyphMix mechanism, which will be introduced later. Subsequently, the mixed image $\mathbf{X}_{g}$ serves as input to train the student model, under the supervision of mixed pseudo-labels $\mathbf{Y}_{g}=\texttt{A}(\mathbf{Y}_{u},\mathbf{Y}_{l})$, by minimizing loss in the following:
@@
To filter out the unreliable predictions in mixed pseudo-labels $\mathbf{Y}_{g}$, we adaptively adjust the matrix $\mathbf{E}$ during pre-training. Specifically, for the labeled synthetic image $\mathbf{X}_{l}$, the positive regions $\mathbf{E}_{l}$ are an all-one matrix. For the unlabeled real image $\mathbf{X}_{u}$, we compute the pixel-wise entropy map on $\mathbf{\hat{Y}}_{u}$ and then apply a dynamic entropy threshold $\zeta$ to obtain its positive regions $\mathbf{E}_{u}$ as follows:
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function. The matrix is subsequently formulated as $\mathbf{E}=\texttt{A}(\mathbf{E}_{u},\mathbf{E}_{l})$. In our experiments, the threshold $\zeta$ is determined by the value corresponding to the top $\gamma$ percent of the entropy map. As the reliability of pseudo-label gradually improves, we smoothly update $\gamma$ using np.linspace(80, 20, t), where t denotes the total number of training iterations.
@@ 3.1 GlyphMix
Although some efforts [63, 60, 30] attempt to aggregate the advantages of synthetic and unlabeled images through versatile transformations, they inevitably introduce intermediate domains, leading to real domain drift. In this section, we present a domain-bridging way tailored for text images, allowing for the free use of synthetic labels while ensuring consistent alignment with the real domain.
1) Graffiti-like Inter-domain Mixing. Given an input synthetic image $\mathbf{X}_{l}$, its corresponding bounding box annotations $\mathcal{B}_{l}$ and converted pixel-wise binary map $\mathbf{Y}_{l}$, our goal is to perform character segmentation. This process identifies all character structure regions and produces a mask $\mathbf{M}_{G}$.
@@
Given the pseudo label $\mathbf{Y}_{u}$ produced by the teacher model on an unlabeled real image $\mathbf{X}_{u}$, we use $\mathbf{M}_{G}$ to generate the inter-domain mixing image and its corresponding label. This process can be formulated as follows:
@@
where 1 is an all-one matrix and $\odot$ denotes the element-wise multiplication operation. $\mathbf{M}_{G}$ indicates pixels occupied by the glyph need to be copied from the synth domain and pasted to the real domain. Thus the mixed image closely aligns with the real domain, taking advantage of the freely available synthetic labels. Put simply, this process produces real-world images with partial annotations derived from synthetic labels.
2) Text-balanced Intra-domain Mixing. Real-world text images inherently exhibit complexity, and the collected unlabeled data often struggles to comprehensively represent the full spectrum of real-world scenarios. Within a student-teacher framework, intra-domain mixing serves as an effective perturbation technique, enriching the diversity of real data and enhancing feature consistency and robustness. However, when randomly generating and applying binary masks to construct intra-domain mixing images and their corresponding labels, there is a risk of missing many text regions, especially considering that text regions typically occupy only a small fraction of an image. To mitigate this challenge, our objective is to retain more text regions while still maintaining randomness.
@@
where 1 is an all-one matrix and $\odot$ represents the element-wise multiplication.